OpenStudy (bashiriz):

URGENT PLEASE HELP: YOU DONT HAVE TO SHOW WORK IF YOU DONT WONT TO I ONLY NEED THE ANSWER! Find the number of permutations in the word “freezer”

OpenStudy (mathmate):

Permutations with repetitions (i.e. identical items) are given by P(n)=n!/(n1!n2!n3!...nk!) when there are k different items, with ni equal to the quantity of each, and n=n1+n2+n3+...+nk. Example: How many ways to arrange 2 apples and 3 oranges on a line? P(5)=5!/(2!3!)=5*4/(1*2)=10 AAOOO AOAOO AOOAO AOOOA OAAOO OAOAO OAOOA OOAAO OOAOA OOOAA